# Test-driven Reinforcement Learning in Continuous Control

### v1

@@ 3.1 Maximum Entropy Reinforcement Learning
Maximum Entropy Reinforcement Learning (MERL, sac_2018 sac_2018) extends the standard reinforcement learning framework by incorporating an entropy maximization objective. This approach simultaneously optimizes for both cumulative reward maximization and policy entropy, thereby improving exploration capacity and algorithmic robustness. MERL extends the standard RL objective by incorporating an entropy regularization term:
@@ 4 Method
In this section, we introduce Test-driven reinforcement learning. We first illustrate the notation and goal in TdRL, then we establish sufficient conditions for trajectory return functions to guarantee policy convergence to the optimal policy set. Additionally, we introduce a lexicographic heuristic approach to compare trajectories for learning trajectory return function. Finally, we detail the implementation of the TdRL algorithm.
4.1 Overview
Figure 1 demonstrates the main procedure of the TdRL algorithm. It follows an iterative procedure mainly consisting of four stages:
@@ 4.1 Overview
Figure 1: The main procedure of the TdRL algorithm, containing trajectory collection, return learning, reward learning and policy learning
@@ 4.4 Lexicographic Trajectory Comparison
For trajectories $\tau_{1}$ and $\tau_{2}$, we need to compare their distances to the optimal trajectory set $\tilde{\mathcal{T}}$, i.e., $\tilde{d}(\tau_{1})$ and $\tilde{d}(\tau_{2})$. However, directly computing $\tilde{d}(\tau)$ is infeasible as $\tilde{\mathcal{T}}$ is unknown. Under such limited information conditions, we employ a lexicographic heuristic (gigerenzer_1996; lexicographic_2020) approach to compare the distances. The lexicographic heuristic is efficient, practical, and adaptable, which is a fast and frugal strategy originally from human decision-making. Specifically, the lexicographic method sequentially evaluates information in priority order and makes decisions based on the first criterion that meets predefined thresholds (e.g., exceeding certain values).
@@ 4.5 TdRL Algorithm
3: Reset $\tau=\emptyset$
4:for each iteration do
5:  Sample action $a$ from $\pi_{\theta}(\cdot|s)$
6:  Exec action $a$ in environment and get $(s^{\prime},done)$
7:  Update Trajectory $\tau\leftarrow\tau\cup\{(s,a,s^{\prime})\}$
8:  Store transition $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s,a,s^{\prime},r_{\phi}(s,a),done)\}$
9:  if donedone then
10:   Store trajectory $\mathcal{D}\leftarrow\mathcal{D}\cup\tau$, reset $\tau=\emptyset$
11:  end if
12:  if iteration $\%K==0$ then
13:   for each gradient step do
14:    Sample minibatch $\{(\tau_{1},\tau_{2})_{j}\}_{j=1}^{D}\sim\mathcal{D}$
15:    Compute $\mu$ for each pair $(\tau_{1},\tau_{2})$ by the lexicographic trajectory comparison approach
16:    Optimize $\mathcal{L}_{R}^{Dis}$ in (16) and $\mathcal{L}_{R}^{Penalty}$ in (17) with respect to $\xi$
17:    Optimize $\mathcal{L}_{r}$ in (18) with respect to $\phi$
18:   end for
19:   Relabel entire replay buffer $\mathcal{B}$ using $r_{\phi}$
20:  end if
21:  for each gradient step do
22:   Sample random minibatch from $\mathcal{B}$
23:   Update policy network $\theta$

### v2

@@ 3.1 Maximum Entropy Reinforcement Learning
Maximum Entropy Reinforcement Learning (MERL, sac_2018 sac_2018) extends the standard reinforcement learning framework by incorporating an entropy maximization objective. This approach simultaneously optimizes for both cumulative reward maximization and policy entropy, thereby improving exploration capacity and algorithmic robustness.The objective of MERL could be written as:
@@ 4 Method
In this section, we introduce Test-driven reinforcement learning. We first illustrate the notation and goal in TdRL, then we establish sufficient conditions for trajectory return functions to guarantee policy convergence to the optimal policy set. Additionally, we introduce a lexicographic heuristic approach to compare trajectories for learning trajectory return function. Finally, we detail the implementation of the TdRL algorithm. Figure 1 demonstrates the main procedure of the TdRL algorithm. It follows an iterative procedure mainly consisting of four stages:
@@
Figure 1: The main procedure of the TdRL algorithm.
@@ 4.3 Lexicographic Trajectory Comparison
For trajectories $\tau_{1}$ and $\tau_{2}$, we need to compare their distances to the optimal trajectory set $\tilde{\mathcal{T}}$, i.e., $\tilde{d}(\tau_{1})$ and $\tilde{d}(\tau_{2})$. However, directly computing $\tilde{d}(\tau)$ is infeasible as $\tilde{\mathcal{T}}$ is unknown. Under such limited information conditions, we employ a lexicographic heuristic (lexicographic_2020) approach to compare the distances. The lexicographic heuristic is efficient, practical, and adaptable, which is a fast and frugal strategy originally from human decision-making (gigerenzer_1996; kk_2013; kk_2021). Specifically, the lexicographic method sequentially evaluates information in priority order and makes decisions based on the first criterion that meets predefined thresholds (e.g., exceeding certain values).
@@ 4.4 TdRL Algorithm
3:for each iteration do
4:  Sample action $a$ from $\pi_{\theta}(\cdot|s)$
5:  Exec action $a$ in environment and get $(s^{\prime},done)$
6:  Update Trajectory $\tau\leftarrow\tau\cup\{(s,a,s^{\prime})\}$
7:  Store transition $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s,a,s^{\prime},r_{\phi}(s,a),done)\}$
8:  if donedone then
9:   Store trajectory $\mathcal{D}\leftarrow\mathcal{D}\cup\tau$, reset $\tau=\emptyset$
10:  end if
11:  if iteration $\%K==0$ then
12:   for each gradient step do
13:    Sample minibatch $\{(\tau_{1},\tau_{2})_{j}\}_{j=1}^{D}\sim\mathcal{D}$
14:    Compute $\mu$ for each pair $(\tau_{1},\tau_{2})$ by the lexicographic trajectory comparison approach
15:    Optimize $\mathcal{L}_{R}^{Dis}$ in (16) and $\mathcal{L}_{R}^{Penalty}$ in (17) with respect to $\xi$
16:    Optimize $\mathcal{L}_{r}$ in (18) with respect to $\phi$
17:   end for
18:   Relabel entire replay buffer $\mathcal{B}$ using $r_{\phi}$
19:  end if
20:  for each gradient step do
21:   Sample random minibatch from $\mathcal{B}$
22:   Update policy network $\theta$
23:  end for
@@ 6 Conclusion
7 Acknowledgments
We thank all anonymous reviewers for their constructive feedback, which helped improve our paper. We also thank Mr. Yuxin Cheng from the University of Hong Kong for his helpful discussions.